# Temporally-Consistent Koopman Autoencoders for Forecasting Dynamical Systems

### v2

@@ Abstract
Absence of sufficiently high-quality data often poses a key challenge in data-driven modeling of high-dimensional spatio-temporal dynamical systems. Koopman Autoencoders (KAEs) harness the expressivity of deep neural networks (DNNs), the dimension reduction capabilities of autoencoders, and the spectral properties of the Koopman operator to learn a reduced-order feature space with simpler, linear dynamics. However, the effectiveness of KAEs is hindered by limited and noisy training datasets, leading to poor generalizability. To address this, we introduce the Temporally-Consistent Koopman Autoencoder (tcKAE), designed to generate accurate long-term predictions even with limited and noisy training data. This is achieved through a consistency regularization term that enforces prediction coherence across different time steps, thus enhancing the robustness and generalizability of tcKAE over existing models. We provide analytical justification for this approach based on Koopman spectral theory and empirically demonstrate tcKAE’s superior performance over state-of-the-art KAE models across a variety of test cases, including simple pendulum oscillations, kinetic plasma, and fluid flow data.
@@ Introduction
This manuscript proposes a new algorithm for consistent long-term prediction of time-series data generated from a nonlinear (possibly high-dimensional) dynamical system using a temporal consistency constraint with Koopman autoencoder framework. This algorithm, dubbed as temporally consistent Koopman autoencoder (tcKAE), compares the predictions from different initial time-instances to a final time in the latent space. This is in contrast with prior KAE methodologies, where the enforcement of multi-step look-ahead prediction loss is reliant on labeled data and/or the latent space representation of the labeled data [6, 7, 8]. Typically larger look-ahead step during training results in greater prediction accuracy, at the cost of training time. However, having large look-ahead step is a constraint when such data is scarce. The proposed tcKAE is different since it evaluates predictions relative to each other, bypassing direct comparison with the encoded version of the labeled data. This alleviates limitations on the maximum look-ahead step imposed by a limited training dataset. However, as discussed in Methods section, selecting arbitrarily large look-ahead steps to enforce consistency among predictions is neither practical nor optimal. Instead, a relatively small number of look-ahead steps can yield comparable or even superior results compared to state-of-the-art Koopman neural network models. The key advantage of the tcKAE algorithm over cKAE [6] lies in its ability to operate without assuming the existence of backward dynamics. It is analytically shown that a KAE spanning a Koopman invariant subspace must satisfy the temporal consistency constraint and by enforcing it, tcKAE leads to higher expressivity and generalizability. This enforcement of temporal consistency in effect reduces the sensitivity of the KAE to noise, decreases the variance or uncertainty in future predictions, and make it more robust with less training data as shown in the experiments. Consistency regularization have been used for semi-supervised learning in classification problems [9, 10, 11], but to the best of our knowledge, this is the first attempt to regularize multi-step temporal consistency for dynamical system learning.
@@
Results
Baseline
(a) $N_{\text{train}}=32$, Clean.
(b) $N_{\text{train}}=32$, 30 dB SNR.
(c) $N_{\text{train}}=64$, Clean.
(d) $N_{\text{train}}=64$, 30 dB SNR.
Performance on Benchmark Datasets
In order to streamline the discussion, we adopt the following notation: the complete dataset is denoted by $\mathbf{X}\in\mathbb{R}^{N_{\text{dim}}\times N_{\text{tot}}}$, where $N_{\text{dim}}$ is the dimension of the state-space, and $N_{\text{tot}}$ is the total number of time samples  [^1] Note that we distinguish between time steps and time samples. By time step, we refer to the discrete time instances used by high-fidelity numerical solvers to generate data. This high-fidelity data is typically sampled at specific time steps and used for KAE training.. We denote the training set as $\mathbf{X}_{\text{train}}\in\mathbb{R}^{N_{\text{dim}}\times N_{\text{train}}}$, validation set as $\mathbf{X}_{\text{val}}\in\mathbb{R}^{N_{\text{dim}}\times N_{\text{val}}}$, and testing set as $\mathbf{X}_{\text{test}}\in\mathbb{R}^{N_{\text{dim}}\times N_{\text{test}}}$ with $N_{\text{train}}$, $N_{\text{val}}$ and $N_{\text{test}}$ being the number of time samples in training, validation and testing set respectively. Note that validation for long-term time-series prediction is nontrivial, and authors in [6] does not mention any validation set. However, for practical applications, validation is crucial, and we validate using the following method: We divide the validation set in two sets $\mathbf{X}_{\text{val}_{1}}\in\mathbb{R}^{N_{\text{dim}}\times N_{\text{val}_{1}}}$ and $\mathbf{X}_{\text{val}_{2}}\in\mathbb{R}^{N_{\text{dim}}\times N_{\text{val}_{2}}}$, separated by $\mathbf{X}_{\text{train}}$ as shown in Figure 1, where $N_{\text{val}}=N_{\text{val}_{1}}+N_{\text{val}_{2}}$. We chose $N_{\text{val}_{1}}=N_{\text{val}_{2}}=\frac{1}{4}N_{\text{train}}$ for all the datasets. The datapoints in $\mathbf{X}_{\text{val}_{1}}$ are used as initial conditions to predict over the span of $\mathbf{X}_{\text{val}_{2}}$. The error metric we use in this work is the 2-norm relative error defined (at $n^{\text{th}}$ time sample) by $\delta_{n}=\frac{||\hat{\mathbf{x}}_{n}-\mathbf{x}_{n}||_{2}}{||\mathbf{x}_{n}||_{2}}$, where $\mathbf{x}$ is the original state and $\hat{\mathbf{x}}$ is the predicted state. The validation loss is determined by calculating the average relative 2-norm error across $\mathbf{X}_{\text{val}_{2}}$, where predictions are generated using initial conditions from points in $\mathbf{X}_{\text{val}_{1}}$. The test error is calculated as the 2-norm relative error in the extrapolation region. To generate an error bar, we use the variance quantified by the 90% confidence interval, defined by the 5th and 95th percentiles of the error. This provides a robust summary of the typical error range by excluding the extreme 5% values at both ends of the distribution. For this statistical analysis, we use $N_{s}$ samples, typically same as $N_{\text{train}}$. The reported results represent the average of neural network training outcomes across 10 different random seeds.
(a) $N_{\text{train}}=28$, Clean.
(b) $N_{\text{train}}=28$, 30 dB SNR.
(c) $N_{\text{train}}=52$, Clean.
(d) $N_{\text{train}}=52$, 30 dB SNR.
Undamped Pendulum
$N_{\text{train}}=32$ $N_{\text{train}}=64$
Method Clean 30 dB SNR Clean 30 dB SNR
DAE 34.411 (20.12) 35.055 (18.83) 13.884 (12.10) 12.421 (9.75)
cKAE 26.415 (18.52) 23.202 (15.71) 12.388 (11.82) 11.316 (9.75)
tcKAE 12.561 (10.20) 15.147 (10.92) 7.491 (6.36) 9.826 (7.91)
Oscillating electron beam
For the electron beam oscillation, one period approximately comprises of 25 time samples. In order to assess the performance of the Koopman models for limited dataset, we train the models for $N_{\text{train}}=28$ and $52$, each for clean and noisy (30 dB SNR) data. For $N_{\text{train}}=28$; $N_{\text{val}_{1}}=N_{\text{val}_{2}}=7$, and for $N_{\text{train}}=52$; $N_{\text{val}_{1}}=N_{\text{val}_{2}}=13$. We extrapolate up to 1000 time samples for evaluating the test error. The snapshots of the electric field distribution and corresponding spatial error profiles for long-term prediction is shown in Figure 4. Note that the spatial error is calculated as following: $\delta(x,y)=\frac{|\mathbf{x}(x,y)-\hat{\mathbf{x}}(x,y)|}{||\mathbf{x}||_{2}}$. As can be seen from Table 2, the advantage of tcKAE is prominent for limited dataset i.e. $N_{\text{train}}=28$.
Table 2: Error (%) comparison for different $N_{\text{train}}$ and noise levels for oscillating electron beam data.
$N_{\text{train}}=28$ $N_{\text{train}}=52$
Method Clean 30 dB SNR Clean 30 dB SNR
DAE 16.570 (4.22) 27.006 (6.20) 3.956 (0.87) 4.782 (0.74)
cKAE 14.958 (2.15) 21.901 (8.28) 3.370 (0.61) 3.900 (0.64)
tcKAE 4.003 (0.73) 6.575 (1.23) 1.489 (0.46) 3.376 (0.52)
(a) $N_{\text{train}}=28$, Clean.
(b) $N_{\text{train}}=28$, 30 dB SNR.
(c) $N_{\text{train}}=40$, Clean.
(d) $N_{\text{train}}=40$, 30 dB SNR.
Flow past cylinder
Flow past cylinder is another common dynamical system used for benchmarking data-driven models. The simulation was carried out using MATLAB’s FEAtool (Figure 8b). A cylinder with diameter 0.1 m is located at a height of 0.2 m. The fluid is characterized by its density $\rho=1$ Kg/m3, and dynamic viscosity $\mu=0.001$ Kg/m s. The flow is unsteady with a maximum velocity of 1 m/s and mean velocity $\frac{2}{3}$ of the maximum velocity. The simulation is run for 80 s until the steady state is achieved, and the horizontal $u$ component of the velocity is probed at every 0.02 s starting from 20 s. The complete dataset $\mathbf{X}\in\mathbb{R}^{2647\times 3001}$ consists of 3001 time samples. More details regarding the simulation setup is presented in Methods: Simulation Setup section. The snapshots and corresponding error profile for long-term prediction is shown in Figure 6. Here also we observe that tcKAE provides significant advantage for limited training data.
Table 3: Error (%) comparison for different $N_{\text{train}}$ and noise levels for flow past cylinder data.
$N_{\text{train}}=28$ $N_{\text{train}}=40$
Method Clean 30 dB SNR Clean 30 dB SNR
DAE 10.182 (2.24) 20.724 (3.60) 1.695 (0.18) 10.775 (1.29)
cKAE 7.153 (1.06) 18.376 (5.58) 1.616 (0.17) 10.933 (1.24)
tcKAE 3.808 (0.45) 13.311 (0.53) 0.874 (0.11) 9.275 (0.94)
In case of flow past cylinder, since one cycle approximately contains 29 time samples, we consider $N_{\text{train}}=28$ and $40$. For $N_{\mathrm{train}}=28$, we take $N_{\text{val}_{1}}=N_{\text{val}_{2}}=7$, and for $N_{\mathrm{train}}=40$, $N_{\text{val}_{1}}=N_{\text{val}_{2}}=10$ is taken. We evaluate the test error for 2500 time samples in the extrapolation region.
Discussion
The proposed tcKAE algorithm advances the state-of-the-art by enforcing multi-step temporal consistency in the latent space. Unlike prior KAE architectures such as consistent Koopman autoencoders (cKAE) [6], which rely on backward dynamics assumptions, tcKAE sidesteps this limitation by introducing a consistency regularization term in the forward direction. The ablation study in the supplementary material further demonstrates that, even without the backward and consistency losses, the combination of forward, identity, and temporal consistency losses achieves comparable or even superior accuracy to cKAE. This innovation strengthens the Koopman invariance of the latent space, enhancing robustness to limited date and noise, improving generalizability. Furthermore, empirical results demonstrate that tcKAE outperforms both cKAE and dynamic autoencoders (DAEs) [7] in terms of long-term prediction accuracy and stability, particularly for scenarios where we have limited training data. The proposed tcKAE not only achieved lower error margins in benchmark problems such as pendulum oscillation, oscillating electron beam, and flow past a cylinder, but also demonstrated narrower confidence intervals, thereby inspiring greater confidence in the predictions.
@@ Methods
where $||\cdot||_{2}$ is the 2-norm, $M$ is the number of samples over which we want to enforce the loss, and $k_{m}$ is the maximum value of $k$, i.e. the maximum look-ahead step for multi-step training. Recent work in [6] demonstrated that including backward dynamics $(K_{b}\in\mathbb{R}^{N_{l}\times N_{l}})$ leads to better stability resulting in more accurate long-term predictions. The idea is to incorporate a backward loss term $(\mathcal{L}_{\text{bwd}})$ with forward-backward consistency loss $(\mathcal{L}_{\text{con}})$,
@@
where $\mathbf{f}^{k}=\mathbf{f}\circ\mathbf{f}\circ\ldots(k\leavevmode\nobreak\ \text{times})\ldots\circ\mathbf{f}$, with $\circ$ being the composition operator. One key aspect of our approach is that instead of enforcing this consistency in the original state space, we do so in the Koopman invariant subspace (latent space). This helps avoid computation in high-dimensional space, and it provides robustness to noise. We justify enforcing such consistency in the latent space by summarizing the desired property of a consistent autoencoder $\Psi_{e}$ and constructing the architecture of this framework. The autoencoder $\Psi_{e}(\cdot)=[\psi_{1}(\cdot),\ldots,\psi_{N_{l}}(\cdot)]$ defines a vector-valued observable of the state space $\mathcal{M}$ where each $\psi_{i}$ is a scalar-valued function. Let $\mathcal{G}$ be the span of $\{\psi_{1},\ldots,\psi_{l}\}$. In order for the linear recurrent matrix $K$ to asymptotically approximate the projection of the Koopman operator $\mathcal{K}$ on $\mathcal{G}$, the latter must be Koopman invariant, i.e., $\mathcal{K}\mathcal{G}\subsetneq\mathcal{G}$. The Koopman invariance will ensure the linear recurrent structure
@@
The proposed architecture introduces a multi-step temporal consistency loss (not against directly encoded version of labelled data) in the latent space in order to strengthen the Koopman invariance of the latent function space $\mathcal{G}$. In the following, we summarize the desirable temporal consistency of the encoder functions upon which the proposed method is constructed.
@@
Let $\Psi_{e}(\cdot)=[\psi_{1}(\cdot),\ldots,\psi_{N_{l}}(\cdot)]^{T}\in\mathcal{F}^{N_{l}}$ denotes a vector valued function comprised with scalar functions $\psi_{i}(\cdot)\in\mathcal{F}$. The latent function space $\mathcal{G}=\{\psi_{1}(\cdot)\ldots,\psi_{N_{l}}(\cdot)\}$ forms a Koopman invariant subspace with respect to the system dynamics (1) if and only if there exists $K\in\mathbb{R}^{N_{l}\times N_{l}}$ such that
@@
()($\Rightarrow$)( ⇒ ) Suppose $\Psi_{e}(\cdot)$ is defined such that $\text{span}(\mathcal{G})$ is a Koopman-invariant subspace. We prove this part through induction. For any $n\geq 0$, (10) is trivially satisfied for $k=1$ with some matrix $K$. Now suppose that it is satisfied for a fixed $k\in\mathbb{N}$. Then, $\Psi_{e}(\mathbf{x}_{n+k+1})=\Psi_{e}\circ\mathbf{f}(\mathbf{x}_{n+k})=\mathcal{K}\Psi_{e}(\mathbf{x}_{n+k})=\mathcal{K}K^{k}\Psi_{e}(\mathbf{x}_{n})=K^{k}\mathcal{K}\Psi_{e}(\mathbf{x}_{n})=K^{k}\Psi_{e}\circ\mathbf{f}(\mathbf{x}_{n})=K^{k}\Psi_{e}(\mathbf{x}_{n+1})=K^{k+1}\Psi_{e}(\mathbf{x}_{n})$, where the first two steps are by the definition of Koopman operator, next two steps are by the linearity of the same, and the last step is by the induction anchor.
()($\Leftarrow$)( ⇐ ) Suppose $\exists$ $K\in\mathbb{R}^{N_{l}\times N_{l}}$ such that (10) is satisfied for all $n\geq 0$ and $k\geq 1$. Then it must be satisfied for $n=0$ and $k=1$, i.e., $\Psi_{e}(\mathbf{x}_{1})=K\Psi_{e}(\mathbf{x}_{0})$. Since $\mathbf{x}_{0}$ can be anything in $\mathcal{M}$, this yields $\Psi_{e}\circ\mathbf{f}(\mathbf{x})=K\Psi_{e}(\mathbf{x})$ for all $\mathbf{x}\in\mathcal{M}$, i.e., $\Psi_{e}\circ\mathbf{f}(\cdot)=K\Psi_{e}(\cdot)$. Let $g\in\text{span}(\mathcal{G})$ be an observable. Then, $\exists$ $\mathbf{\alpha}\in\mathbb{R}^{l}$ such that $g(\cdot)=\mathbf{\alpha}^{T}\Psi_{e}(\cdot)$. Now, $\mathcal{K}g(\cdot)=g\circ\mathbf{f}(\cdot)=\mathbf{\alpha}^{T}\Psi_{e}\circ\mathbf{f}(\cdot)=\mathbf{\alpha}^{T}K\Psi_{e}(\cdot)$, i.e., $\mathcal{K}g\in\text{span}(\mathcal{G})$. Hence, $\text{span}(\mathcal{G})$ forms a Koopman-invariant subspace. ∎
@@
Figure 9: tcKAE architecture where the layers with red nodes indicate the input and output layer in the original state space. The latent space is represented by the bottleneck layer with blue nodes. The hidden layers are shown by the gray nodes.
@@
The network architecture is shown in Figure 9. Note that the network topology is same as that in [6]. The number of nodes in the input and output layer (red) is indicated by $N_{\text{in}}$ and $N_{\text{out}}$ ($N_{\text{in}}=N_{\text{out}}$). $N_{l}$ denotes the number of nodes in the bottleneck layer (blue), which represents the approximation of Koopman-invariant latent space. Number of nodes in each hidden layers (two hidden layers for each encoder and decoder) is represented by $N_{h}$ (gray nodes). We keep $N_{h}$ same for both the hidden layers and it is tuned in multiple of $16$. We use the $tanh$ activation function for each layer except the bottleneck layer.

### v3

# Temporally Consistent Koopman Autoencoders for Forecasting Dynamical Systems

@@ Abstract
Absence of sufficiently high-quality data often poses a key challenge in data-driven modeling of high-dimensional spatio-temporal dynamical systems. Koopman Autoencoders (KAEs) harness the expressivity of deep neural networks (DNNs), the dimension reduction capabilities of autoencoders, and the spectral properties of the Koopman operator to learn a reduced-order feature space with simpler, linear dynamics. However, the effectiveness of KAEs is hindered by limited and noisy training datasets, leading to poor generalizability. To address this, we introduce the temporally consistent Koopman autoencoder (tcKAE), designed to generate accurate long-term predictions even with limited and noisy training data. This is achieved through a consistency regularization term that enforces prediction coherence across different time steps, thus enhancing the robustness and generalizability of tcKAE over existing models. We provide analytical justification for this approach based on Koopman spectral theory and empirically demonstrate tcKAE’s superior performance over state-of-the-art KAE models across a variety of test cases, including simple pendulum oscillations, kinetic plasma, and fluid flow data.
@@ Introduction
This manuscript proposes a new algorithm for consistent long-term prediction of time-series data generated from a nonlinear (possibly high-dimensional) dynamical system using a temporal consistency constraint with Koopman autoencoder framework. This algorithm, dubbed as temporally-consistent Koopman autoencoder (tcKAE), compares the predictions from different initial time-instances to a final time in the latent space. This is in contrast with prior KAE methodologies, where the enforcement of multi-step look-ahead prediction loss is reliant on labeled data and/or the latent space representation of the labeled data [6, 7, 8]. Typically larger look-ahead step during training results in greater prediction accuracy, at the cost of training time. However, having large look-ahead step is a constraint when such data is scarce. The proposed tcKAE is different since it evaluates predictions relative to each other, bypassing direct comparison with the encoded version of the labeled data. This alleviates limitations on the maximum look-ahead step imposed by a limited training dataset. However, as discussed in Methods section, selecting arbitrarily large look-ahead steps to enforce consistency among predictions is neither practical nor optimal. Instead, a relatively small number of look-ahead steps can yield comparable or even superior results compared to state-of-the-art Koopman neural network models. The key advantage of the tcKAE algorithm over cKAE [6] lies in its ability to operate without assuming the existence of backward dynamics. It is analytically shown that a KAE spanning a Koopman invariant subspace must satisfy the temporal consistency constraint and by enforcing it, tcKAE leads to higher expressivity and generalizability. This enforcement of temporal consistency in effect reduces the sensitivity of the KAE to noise, decreases the variance or uncertainty in future predictions, and make it more robust with less training data as shown in the experiments. Consistency regularization have been used for semi-supervised learning in classification problems [9, 10, 11], but to the best of our knowledge, this is the first attempt to regularize multi-step temporal consistency for dynamical system learning.
@@ Methods
Figure 1: cKAE architecture where the layers with red nodes indicate the input and output layer in the original state space. The latent space is represented by the bottleneck layer with blue nodes. The hidden layers are shown by the gray nodes. The encoder and decoder layers are denoted by $\Psi_{e}$ and $\Psi_{d}$ respectively. The linear layers represent the linear forward $(K)$ and backward dynamics $(K_{b})$.
@@
where $||\cdot||_{2}$ is the 2-norm, $M$ is the number of samples over which we want to enforce the loss, and $k_{m}$ is the maximum value of $k$, i.e. the maximum look-ahead step for multi-step training. Recent work in [6] demonstrated that including backward dynamics $(K_{b}\in\mathbb{R}^{N_{l}\times N_{l}},\leavevmode\nobreak\ \Psi_{d}\circ K_{b}^{k}\circ\Psi_{e}(\mathbf{x}_{n})=\hat{\mathbf{x}}_{n-k})$ as shown in Figure 1 leads to better stability resulting in more accurate long-term predictions. The idea is to incorporate a backward loss term $(\mathcal{L}_{\text{bwd}})$ with forward-backward consistency loss $(\mathcal{L}_{\text{con}})$,
@@
where $\mathbf{f}^{k}=\mathbf{f}\circ\mathbf{f}\circ\ldots(k\leavevmode\nobreak\ \text{times})\ldots\circ\mathbf{f}$, with $\circ$ being the composition operator. One key aspect of our approach is that instead of enforcing this consistency in the original state space, we do so in the Koopman invariant subspace (latent space). This helps to avoid computation in high-dimensional space and provides robustness to noise. We justify enforcing such consistency in the latent space by summarizing the desired property of a consistent autoencoder $\Psi_{e}$ and constructing the architecture of this framework. The autoencoder $\Psi_{e}(\cdot)=[\psi_{1}(\cdot),\ldots,\psi_{N_{l}}(\cdot)]$ defines a vector-valued observable of the state space $\mathcal{M}$ where each $\psi_{i}$ is a scalar-valued function. Let $\mathcal{G}$ be the span of $\{\psi_{1},\ldots,\psi_{N_{l}}\}$. In order for the linear recurrent matrix $K$ to asymptotically approximate the projection of the Koopman operator $\mathcal{K}$ on $\mathcal{G}$, the latter must be Koopman invariant, i.e., $\mathcal{K}\mathcal{G}\subsetneq\mathcal{G}$. The Koopman invariance will ensure the linear recurrent structure
@@
The proposed architecture introduces a multi-step temporal consistency loss (not applied against the directly encoded version of the labeled data) in the latent space in order to strengthen the Koopman invariance of the latent function space $\mathcal{G}$. It is important to emphasize that, for calculating temporal consistency, comparisons are made between multiple forward predictions in the latent space. These predictions are not necessarily constrained by the encoded versions of the available training data. In the following, we summarize the desirable temporal consistency of the encoder functions upon which the proposed method is constructed.
@@
Let $\Psi_{e}(\cdot)=[\psi_{1}(\cdot),\ldots,\psi_{N_{l}}(\cdot)]^{T}\in\mathcal{F}^{N_{l}}$ denote a vector valued function (encoder) comprised of scalar functions $\psi_{i}(\cdot)\in\mathcal{F}$. The latent space $\mathcal{G}=\text{span}\{\psi_{1}(\cdot)\ldots,\psi_{N_{l}}(\cdot)\}$ forms a Koopman invariant subspace with respect to the system dynamics (1) if and only if there exists $K\in\mathbb{R}^{N_{l}\times N_{l}}$ such that
@@
()($\Rightarrow$)( ⇒ ) Suppose $\Psi_{e}(\cdot)$ is defined such that $\mathcal{G}$ is a Koopman-invariant subspace. We prove this part through induction. For any $n\geq 0$, (10) is satisfied for $k=1$ with a matrix $K$ where $K=\mathcal{K}_{\mathcal{G}}$, i.e., the restriction of the Koopman operator $\mathcal{K}$ on $\mathcal{G}$. Now suppose that it is satisfied for a fixed $k\in\mathbb{N}$. Then, $\Psi_{e}(\mathbf{x}_{n+k+1})=\Psi_{e}\circ\mathbf{f}(\mathbf{x}_{n+k})=\mathcal{K}\Psi_{e}(\mathbf{x}_{n+k})=\mathcal{K}K^{k}\Psi_{e}(\mathbf{x}_{n})=K^{k}\mathcal{K}\Psi_{e}(\mathbf{x}_{n})=K^{k}\Psi_{e}\circ\mathbf{f}(\mathbf{x}_{n})=K^{k}\Psi_{e}(\mathbf{x}_{n+1})=K^{k+1}\Psi_{e}(\mathbf{x}_{n})$, where the first two steps are by the definition of Koopman operator, next two steps are by the linearity of the same, and the last step is by the induction anchor.
()($\Leftarrow$)( ⇐ ) Suppose $\exists$ $K\in\mathbb{R}^{N_{l}\times N_{l}}$ such that (10) is satisfied for all $n\geq 0$ and $k\geq 1$. Then it must be satisfied for $n=0$ and $k=1$, i.e., $\Psi_{e}(\mathbf{x}_{1})=K\Psi_{e}(\mathbf{x}_{0})$. Since $\mathbf{x}_{0}$ can be anything in $\mathcal{M}$, this yields $\Psi_{e}\circ\mathbf{f}(\mathbf{x})=K\Psi_{e}(\mathbf{x})$ for all $\mathbf{x}\in\mathcal{M}$, i.e., $\Psi_{e}\circ\mathbf{f}(\cdot)=K\Psi_{e}(\cdot)$. Let $g\in\mathcal{G}$ be an observable. Then, $\exists$ $\mathbf{\alpha}=[\alpha_{1},\ldots,\alpha_{N_{l}}]\in\mathbb{R}^{N_{l}}$ such that $g(\cdot)=\sum\limits_{i=1}^{N_{l}}\alpha_{i}\psi_{N_{i}}(\cdot)=\mathbf{\alpha}^{\top}\Psi_{e}(\cdot)$. Now, $\mathcal{K}g(\cdot)=g\circ\mathbf{f}(\cdot)=\mathbf{\alpha}^{\top}\Psi_{e}\circ\mathbf{f}(\cdot)=\mathbf{\alpha}^{\top}K\Psi_{e}(\cdot)=\sum\limits_{j=1}^{N_{l}}\left(\sum\limits_{i=1}^{N_{l}}\alpha_{i}K_{ij}\right)\psi_{j}(\cdot)$, i.e., $\mathcal{K}g\in\mathcal{G}$. Hence, $\mathcal{G}=\text{span}\{\psi_{1}(\cdot)\ldots,\psi_{N_{l}}(\cdot)\}$ forms a Koopman-invariant subspace. ∎
@@
The overall KAE network architecture is shown in Figure 1 and the tcKAE architecture is shown in Figure 2. Note that the network topology is same as that in [6]. The number of nodes in the input and output layer (red) is indicated by $N_{\text{in}}$ and $N_{\text{out}}$ ($N_{\text{in}}=N_{\text{out}}$). $N_{l}$ denotes the number of nodes in the bottleneck layer (blue), which represents the approximation of Koopman-invariant latent space. Number of nodes in each hidden layers (two hidden layers for each encoder and decoder) is represented by $N_{h}$ (gray nodes). We keep $N_{h}$ same for both the hidden layers and it is tuned in multiple of $16$. We use the $tanh$ activation function for each layer except the bottleneck layer.